OpenStudy (moonlitfate):

Find the exact length of the curve x= 9 cos t - cos 9t , y = 9 sin t - sin 9t , 0

OpenStudy (moonlitfate):

This is what I have so far: \[L = \int\limits_{0}^{\pi}\sqrt{(9 \cos t -9 \cos 9t)^2 + (-9 \sin t + 9 \sin9t)^2}dt\]

OpenStudy (moonlitfate):

Then you expand that out to be \[\int\limits_{0}^{\pi}\sqrt{81 \cos^2t-162 \cos t \cos 9t + 81 \cos^2 9t + 81 \sin^2 t - 162 \sin t \sin 9t + 81 \sin^2 9t} dt\]

OpenStudy (irishboy123):

take the 9 outside the root!

OpenStudy (moonlitfate):

The issue that I'm having is just simplifying everything.

OpenStudy (irishboy123):

1/ combine the cos^2 + sin^2 pairs

OpenStudy (moonlitfate):

Would that also include the sin^2(9t) and cos^2(9t) as well. I apologize if that seems like a silly question. I just wasn't sure.

OpenStudy (irishboy123):

yes! but please get those 81's and 162'2 out and in fornt of the \(\int\)!!! it's trivial but it doesn't help

OpenStudy (moonlitfate):

Okay! Just give me a moment.:) Working this out as I go, and I thank you for your help and patience so far.

OpenStudy (irishboy123):

take your time, there is no rush

OpenStudy (moonlitfate):

It's easier for me to group these. Way to easy to get a bit thrown off by all the stuff going on under the radical. \[9 \sqrt{9 (1) + 9(1) -9 [2 (\cos t \cos 9t + 9 \sin t \sin 9t)}\]

OpenStudy (moonlitfate):

Hopefully, I didn't make a mistake somewhere.

OpenStudy (irishboy123):

the 9 should go out completely. you know you can copy your latex and re-do it, instead or re-typing it all?

OpenStudy (moonlitfate):

I definitely forget that I can do that.

OpenStudy (irishboy123):

double tap or right click on the equation and one of the options will be to show the latex code this will save you so much time in the long run if you can do that i use chrome then you can paste it back inside \( or \[ and \) or \]

OpenStudy (irishboy123):

https://gyazo.com/f991fb8f9a929c36365c5c8843d3543d take show TEX then Ctrl A and Ctrl C!!

OpenStudy (moonlitfate):

Let me see if I actually have the correct simplification of this radical. So far. \[9 \sqrt{(1 + 1 ) -2 (\cos t \cos 9t + \sin t \sin 9t)}\] \[9 \sqrt{(2 ) -2 (\cos t \cos 9t + \sin t \sin 9t)}\]

OpenStudy (irishboy123):

\[3 \sqrt{2} \sqrt{1 - (\cos t \cos 9t + \sin t \sin 9t)}\]

OpenStudy (moonlitfate):

Always that algebra messing me up, but that's okay. At least I know I'm on the right track.

OpenStudy (irishboy123):

do you agree with the algebra? need my own work checked, you see :-)

OpenStudy (moonlitfate):

It looks fine me.

OpenStudy (moonlitfate):

So, the part that confuses me is that trig piece.

OpenStudy (irishboy123):

nah, it's: \[9 \sqrt{2} \sqrt{1 - (\cos t \cos 9t + \sin t \sin 9t)}\]

OpenStudy (irishboy123):

we can simplify: \[(\cos t \cos 9t + \sin t \sin 9t)\]

OpenStudy (irishboy123):

hint double angle formulae?? for cosine??

OpenStudy (moonlitfate):

I have to look that up real quick. ;/ Something I should know off the top of my head at this point.

OpenStudy (irishboy123):

no rush!

OpenStudy (moonlitfate):

Double angle for cosine: \[\cos (2u) = \cos^2 u - \sin ^2 u = 2 \cos^2 u - 1= 1 - 2 \sin^2 u \]

OpenStudy (irishboy123):

yes last one looks good ;-)

OpenStudy (moonlitfate):

Trying to get that fit into the double angle is just what's confusing me. I'm sorry.

OpenStudy (irishboy123):

sorry, my bad, i was looking a step ahead let's start with \(cos (A- B) = cosA cosB + sinA sinB\) can you work that and \((\cos t \cos 9t + \sin t \sin 9t)\)

OpenStudy (moonlitfate):

\[\cos (t- 9t) = costcos9t + sintsin9t\]

OpenStudy (irishboy123):

brill and cos (-x) = cos x so next step is?

OpenStudy (moonlitfate):

Would that give you \[\cos (-8t) = - \cos(8t) \]

OpenStudy (irishboy123):

no + cos 8t

OpenStudy (moonlitfate):

Oh, mixed up my signs. I got it.

OpenStudy (irishboy123):

cos(-x) = cos x [going to relaunch this thread as my copy/ paste thing is being funny "not" going away - and this one has a really interesting ending

OpenStudy (moonlitfate):

So, would I have to plug that in to the last part of the double angle formula for cosine and get: \[\cos(8t) = 1- 2 \sin^2(8t) \]

OpenStudy (irishboy123):

yes post it!

OpenStudy (moonlitfate):

So, I would plug that back into the integral with the crazy trig: \[9\sqrt{2} \int\limits_{0}^{\pi}\sqrt{1-(1-2 \sin^2(8t)} dt\]

OpenStudy (irishboy123):

soz again \[\cos(8t) = 1- 2 \sin^2(4t)\] there was a typo there - because we half the angle yes :p

OpenStudy (irishboy123):

take it from here \[9\sqrt{2} \int\limits_{0}^{\pi}\sqrt{1-(1-2 \sin^2(4t)} ) \ dt\]

OpenStudy (irishboy123):

that is what we are measuring! believe it or not. the folks at desmos have asked me if it can be used as some kind of staff favourite. feels like a practical joke!!

OpenStudy (irishboy123):